# Learning to Explore with Parameter-Space Noise: A Deep Dive into Parameter-Space Noise for Reinforcement Learning with Verifiable Rewards

### v2

@@ 1 Introduction
Reinforcement Learning with Verifiable Rewards (RLVR) has become a central paradigm for improving the reasoning behaviors of Large Language Models (LLMs), delivering substantial gains on domains with automatic correctness signals such as mathematics and code generation  (Lu et al., 2024; Ouyang and others, 2022; Lambert et al., 2025). In this setting, policy-gradient style algorithms—most notably Proximal Policy Optimization (PPO)  (schulman2017proximal) and its recent variants such as Group Relative Policy Optimization (GRPO)  (Guo et al., 2025)—enable direct optimization against verifiers (e.g., unit tests or symbolic checkers), aligning models with ground-truth correctness and propelling systems like DeepSeek-R1 (Guo et al., 2025) to strong empirical performance.
@@ 4.2 Results and Q&As
A: Yes. We demonstrate that PSN is model-agnostic by training on Qwen2.5-7B. Furthermore, both theoretical intuition and empirical results identify MLP layers as the optimal injection site for maximizing the reasoning capability boundary.
@@
Theoretical work by Plappert et al. (Plappert et al., 2018) suggests that parameter space noise yields optimal performance when using normalization between perturbed layers. In standard Transformer architectures, the MLP blocks naturally satisfy this structural criterion. To empirically validate this hypothesis and demonstrate the method’s generalization beyond specific math-tuned models, we conducted ablation studies using the general-purpose Qwen2.5-7B model. We compared noise injection across three distinct loci: full layers, the language modeling head (lm_head), and the MLP blocks. As illustrated in Figure 4, injecting noise exclusively into the MLP layers yields a superior reasoning capability boundary, demonstrating significantly higher pass@$k$ performance at large sampling budgets ($k\geq 128$) compared to other strategies. Comprehensive results are provided in Appendix Figure 10.
@@
To probe whether the gains in reasoning capability boundary reflect genuine exploration (rather than mere reweighting), we conduct an Gemini-assisted (Team, 2025a) qualitative analysis on AIME 2024 under our standard evaluation protocol (30 problems; n=300n{=}300 rollouts per problem). We focus on the subset of problems where the base/original model fails on all $300$ rollouts, yet PSN-RLVR produces at least one correct rollout. Across these cases, we find that the successful PSN-RLVR traces typically employ solution perspective that are absent from the base model’s rollout set, indicating that PSN-RLVR can access new reasoning modes rather than only improving selection among pre-existing trajectories. Representative examples are provided in Appendix I.
@@ Appendix B Training experiment setting
Appendix D Detailed performance of training time temperature scaling
@@
Appendix E More experiment result of evaluation temperature scaling
@@
Appendix F More detailed experiment result of evaluation temperature scaling
@@
Appendix G Detailed performance of where to inject noise
@@
Appendix H Detailed performance of adaptive noise
@@
Appendix I Detailed example of PSN-GRPO discovers qualitatively new solution strategies

### v3

# arXiv:2602.02555v3[cs.LG] 28 Feb 2026

@@ 1 Introduction
Reinforcement Learning with Verifiable Rewards (RLVR) has become a central paradigm for improving the reasoning behaviors of Large Language Models (LLMs), delivering substantial gains on domains with automatic correctness signals such as mathematics and code generation  (Lu et al., 2024; Ouyang and others, 2022; Lambert et al., 2025). In this setting, policy-gradient style algorithms—most notably Proximal Policy Optimization (PPO)  (Schulman et al., 2017) and its recent variants such as Group Relative Policy Optimization (GRPO)  (Guo et al., 2025)—enable direct optimization against verifiers (e.g., unit tests or symbolic checkers), aligning models with ground-truth correctness and propelling systems like DeepSeek-R1 (Guo et al., 2025) to strong empirical performance.
@@ 4.2 Results and Q&As
A: Yes. We demonstrate that PSN is model-agnostic by training on Qwen2.5-7B and Qwen3-4B. Furthermore, both theoretical intuition and empirical results identify MLP layers as the optimal injection site for maximizing the reasoning capability boundary.
@@
Theoretical work by Plappert et al. (Plappert et al., 2018) suggests that parameter space noise yields optimal performance when using normalization between perturbed layers. In standard Transformer architectures, the MLP blocks naturally satisfy this structural criterion. To empirically validate this hypothesis and demonstrate the method’s generalization beyond specific math-tuned models, we conducted ablation studies using the general-purpose Qwen2.5-7B model. We compared noise injection across three distinct location: full decoding layers, the language modeling head (lm_head), and the MLP blocks. As illustrated in Figure 4, injecting noise exclusively into the MLP layers yields a superior reasoning capability boundary, demonstrating significantly higher pass@$k$ performance at large sampling budgets ($k\geq 128$) compared to other strategies. Comprehensive results are provided in Appendix Figure 10.
Additional evidence on Qwen3-4B-Base.
[…]
Method (Qwen3-4B-Base) Avg pass@1 Avg pass@2 Avg pass@256
Clean ($\sigma=0$) 33.8% 41.5% 72.0%
PSN ($\sigma=0.001$) 31.9% 40.5% 72.5%
Table 1: Qwen3-4B-Base generalization. Average pass@$k$ (%) over five math benchmarks (AIME24/25, AMC23, Minerva-Math, OlympiadBench) comparing clean decoding vs. PSN. PSN expands the high-budget reasoning boundary (avg pass@256 +0.5pp) while slightly reducing low-budget reliability, matching the exploration–exploitation trade-off.
@@
To probe whether the gains in reasoning capability boundary reflect genuine exploration (rather than mere reweighting), we conduct an Gemini-assisted (Team, 2025a) qualitative analysis on AIME 2024 under our standard evaluation protocol (30 problems; n=300n{=}300 rollouts per problem). We focus on the subset of problems where the base/original model fails on all $300$ rollouts, yet PSN-RLVR produces at least one correct rollout. Across these cases, we find that the successful PSN-RLVR traces typically employ solution perspective that are absent from the base model’s rollout set, indicating that PSN-RLVR can access new reasoning modes rather than only improving selection among pre-existing trajectories. Representative examples are provided in Appendix J.
@@ Appendix B Training experiment setting
Appendix D Result of Qwen3-4B-Base
Benchmark Clean ($\sigma=0$) PSN ($\sigma=0.001$) Gap (PSN-Clean)
pass@1 pass@2 pass@256 pass@1 pass@2 pass@256 pass@1 pass@2 pass@256
AIME 24 19.4% 25.7% 62.8% 17.4% 23.6% 65.5% -2.0 -2.1 +2.7
AIME 25 15.5% 20.6% 58.9% 14.7% 20.0% 62.6% -0.8 -0.6 +3.7
Minerva-Math 32.2% 39.3% 63.6% 32.0% 40.7% 64.8% -0.2 +1.4 +1.2
OlympiadBench 44.4% 52.9% 77.1% 40.6% 50.8% 77.4% -3.8 -2.1 +0.3
Table 7: Qwen3-4B-Base detailed results (clean vs. PSN). Per-benchmark pass@$k$ (%) for Qwen3-4B-Base comparing clean decoding ($\sigma=0$) and PSN ($\sigma=0.001$; MLP-only perturbation). PSN exhibits the typical exploration–exploitation trade-off: slightly lower low-budget reliability (pass@1/2) but an expanded high-budget capability boundary (pass@256), especially on harder AIME benchmarks.
Appendix E Detailed performance of training time temperature scaling
@@
Appendix F More experiment result of evaluation temperature scaling
@@
Appendix G More detailed experiment result of evaluation temperature scaling
@@
Appendix H Detailed performance of where to inject noise
@@
Appendix I Detailed performance of adaptive noise
@@
Appendix J Detailed example of PSN-GRPO discovers qualitatively new solution strategies